OpenStudy (anonymous):

Find an equation of the tangent line to the curve at the given point. y = sin(sin x), (3π, 0)

jimthompson5910 (jim_thompson5910):

what is y' equal to?

OpenStudy (anonymous):

sin(sin(x))

jimthompson5910 (jim_thompson5910):

y is equal to that, you have to derive both sides with respect to x to get y '

jimthompson5910 (jim_thompson5910):

or dy/dx

jimthompson5910 (jim_thompson5910):

use the chain rule

OpenStudy (anonymous):

The thing is I don't understand the chain rule. I'm taking my first online class and my teacher is horrible.

OpenStudy (anonymous):

what is the derivative of sine?

OpenStudy (anonymous):

cosine

OpenStudy (anonymous):

ok right so the chain rule the derivative of the sine of something is the cosine of something, times the derivative of the something for example, \[\frac{d}{dx}[\sin(x^2)]=\cos(x^2)\times 2x\] and \[\frac{d}{dx}\sin(e^x)=\cos(e^x)\times e^x\]

OpenStudy (anonymous):

in your case you have \[\sin(\sin(x))\] so the derivative is \[\cos(\sin(x))\times \cos(x)\]

jimthompson5910 (jim_thompson5910):